Listed below are the top 10 annual salaries​ (in millions of​ dollars) of TV personalities. Find the​ (a) mean,​ (b) median,​ (c) mode, and​ (d) midrange for the given sample data in millions of dollars. Given that these are the top 10​ salaries, do we know anything about the salaries of TV personalities in​ general? Are such top 10 lists valuable for gaining insight into the larger​ population? 37.7    36.8    35.4    27.7    15.4    12.2    12.8    10.2    8.9    8.8
Answer #1

Solution:-

8.8 8.9 10.2 12.2 12.8 15.4 27.7 35.4 36.8 37.7

a.)

Mean=sum of all 10 TV personalities annual salaries/10

=(37.7+36.8+5.4+27.7+15.4+12.2+12.8+10.2+ 8.9+8.8)/10

=20.59

b.)Median=(12.8+15.4)/2

=14.1

c.) There is No mode for the given data.

d.)Mid range=(Lowest value+Highest Value)/2

=(8.8+37.7)/2

=23.25

Next Part:-

Ans:- No, Because here samples are top 10​ salaries, No information is given.

Next part:

ans:-No. because such top 10 lists represent an extreme subset of the population.